OpenStudy (anonymous):

Multiply: (8a + 5b) (8a - 5b). A.16a2 + 25b2 B. 64a2 - 25b2 C.16a2 - 10ab + 25b2 D. 64a2 - 40ab + 25b

OpenStudy (anonymous):

\[64a ^{2 }- 25b ^{2}\]

OpenStudy (anonymous):

64a^2-25b^2

OpenStudy (mayliz):

B..64a2-25b2

OpenStudy (anonymous):

Thanks

OpenStudy (callisto):

Use difference of 2 squares: (a-b)(a+b) = a^2 -b^2 put a = 8a and b= 5b into the formula , you'll get (8a + 5b) (8a - 5b) = (8a)^2 - (5b)^2 = 64a^2 - 25b^2